OpenStudy (anonymous):

Extracting the roots Solve: 9a^2=54

OpenStudy (anonymous):

divide by 9 then take the square root make sure to write two answers, the positive and the negative one

OpenStudy (anonymous):

ishaan i have a question

OpenStudy (anonymous):

whats the answer?

OpenStudy (anonymous):

ok here we go \[9a^2=54\] \[a^2=\frac{54}{9}=6\] \[a=\pm\sqrt{6}\]